write each rational number in the form a/b where a and b are integers and b = 0.
a.) 0.3
b.)5 1/3
c.)0
d.)73
e.)-0.21

Answer by stanbon(48545)

write each rational number in the form a/b where a and b are integers and b = 0.
a.) 0.3 = 3/10
b.)5 1/3 = 16/3
c.)0 = 0/1
d.)73 = 73/1
e.)-0.21 = -21/100
========================
Cheers,
Stan H.